OpenStudy (anonymous):

How to find the derivative of f(x)= 1/(x+1)(2x-3) using the qoutient rule

OpenStudy (anonymous):

don't use it \[\left(\frac{1}{f}\right)'=-\frac{f'}{f^2}\]

OpenStudy (anonymous):

only need the quotient rule if there is a variable in the denominator

OpenStudy (anonymous):

What do you mean with that

OpenStudy (anonymous):

he means let f = (x+1)(2x-3) and use the formula he wrote...

OpenStudy (anonymous):

you wanna just do the regular quotient rule?

OpenStudy (anonymous):

Yes plz...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):